An appliance connected to the 230v mains supply has a resistance of 23 ohms and is on for 2.5 hours per day. Work out how much energy the appliance uses per day and give an appropriate unit
1
Expert's answer
2018-10-23T11:22:09-0400

Voltage:


U=230V

Resistance:

R= 23 ohms

Time:

t= 2.5 hour

Solution:

As you know, work A is equal to:

A=Pt

where P – power:

P=U^2/R=〖230〗^2/23=2300W

A=2300/1000*2.5=5.75kWh

Answer:

Energy the appliance uses per day:

A=5.75kwh

or

A=20.7MJ